OpenStudy (anonymous):

Find a ⋅ b. a = 8i + 6j, b = 4i + 5j

OpenStudy (anonymous):

You are to just mutiply them.. Do you know about FOIL??

OpenStudy (anonymous):

yes but the answer choices are numbers! i know i as an imaginary number but j as a variable and i think they are both variables in the problem.

OpenStudy (anonymous):

If i is an imaginary here then : \[-32 + 64i.j + 30j^2\]

OpenStudy (anonymous):

its not............... the answer is a digit....... idk what it is

OpenStudy (anonymous):

no its somethign to do with dot products and vectors

OpenStudy (anonymous):

\[a \color{red} {\cdot} b = |ab|\cos(\theta)\]

OpenStudy (anonymous):

Or simply mutiply i coefficient with i and j with j.. \[a \cdot b = 8 \times 4 + 6 \times 5 \implies \color{green}{32 + 30} = ??\]

OpenStudy (anonymous):

yes!!! thank you!!!

OpenStudy (anonymous):

Got it right or not?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

now do you know if i am multiplying v *w and v = <3, -8, -3>; w = <-4, -2, -6> how do i solve this?

OpenStudy (anonymous):

like do i multiply v1*w1 + v2w2+v3w3?

OpenStudy (anonymous):

Is it \(v \cdot w\) or \(v \times w\) ??

OpenStudy (anonymous):

v⋅w

OpenStudy (anonymous):

Yes just multiply the corresponding elemets there..

OpenStudy (anonymous):

Yes, as you are saying go like that... v1*w1 + v2......... and so on..

OpenStudy (anonymous):

thank you:)